### v1

@@ 2.1 Long-Context Processing with LLMs
1:$input$ $sequence$$\{I\},$$\#layers$$\{L\},$ $TSPlayer,$ $TSPlength$$\{B_{TSP}\},$ $KVbudget$$\{B_{KV}\}$
@@
5:     if $l\leq TSPlayer$ then
6:         $X,Att_{l},K_{X},V_{X}\leftarrow layer_{l}(X)$
@@
8:         if $l==TSPlayer$ then
9:              $x\leftarrow HiddenCompress(X,Att_{l},B_{TSP})$
@@
2.2 KV Cache Compression
KV cache compression techniques for long-context processing are primarily designed to prune tokens, thereby reducing the associated KV caches and effectively managing memory usage. A common approach leverages the attention scores generated by each attention head to selectively prune tokens in an attention-head-wise manner. This method operates on the premise that token importance varies across attention heads. A notable example is SnapKV Li et al. (2024) in Figure 2(a). It discovers that each attention head in the model consistently focuses on specific prompt token during generation, and this robust pattern can be obtained from an observation window located at the end of the input prompt. Thus, SnapKV evaluates token importance score for each attention head by summing the attention scores assigned to each token as follows:
Here, $S_{i}^{l,h}$ is the importance score of $i$-th token in $h$-th attention head of the $l$-th layer. $Att_{l}$ denotes the attention score matrix of $l$-th layer, while $N_{I}$ and $N_{obs}$ indicate the number of tokens in the input prompt and the observation window, respectively. SnapKV applies average pooling on the scores to prevent sparse token selection. Here, $12w_{p}+$ 1 indicates the size of pooling window. SnapKV selects tokens from the observation window first and then selects tokens with the highest importance scores for KV caching. Since SnapKV relies on the attention scores produced by each attention head, it requires propagationg of the full input prompt throughout the entire model, during the prefill stage. As a result, SnapKV cannot improve computational efficiency or reduce the latency of the prefill stage, which directly impacts TTFT, a key factor in optimizing user experience for LLM serving.
AdaKV Feng et al. (2024) in Figure 2(b) expands SnapKV approach to grouped-query-attention (GQA) mechanism Ainslie et al. (2023) by introducing dynamic KV budget allocation across attention groups and selecting important tokens in attention-group-wise manner. To enhance accuracy preservation, HeadKV Fu et al. (2024) in Figure 2(c) builds upon SnapKV by introducing dynamic KV budget allocation across attention heads. This feature has shown effective in preserving accuracy under aggressive KV cache compression scenarios, such as when the average KV budget is as low as 128. However, when applied to more realistic KV budget targets, such as 512, dynamic KV budget allocation exhibits marginal accuracy improvements or even a slight accuracy degradation compared to SnapKV, while they introduces notable latency and throughput overhead. The primary reason of the overhead is that the processing speed of the attention mechanism is limited by the attention head with the longest KV cache. As the size of KV budget increases, the imbalance in KV cache sizes across attention heads becomes more pronounced and it amplifies latency overhead. This overhead makes these methods less practical for real-world applications, where achieving low latency is critical.
In contrast, GemFilter Shi et al. (2024) in Figure 2(d) adopts a fundamentally different approach that enhances the computational efficiency of the prefill stage. GemFilter is based on the observation that LLMs can identify relevant tokens in intermediate layers, referred to as the filter layer. The filter layer selects the indices of relevant tokens at the layer level rather then head level, and GemFilter uses these indices to compress the input prompt. Then, the prefill stage is recomputed with compressed input prompt. By processing full-context with part of the LLM, GemFilter effectively reduces TTFT. The GemFilter paper reports successful accuracy preservation when using a KV cache budget of 2k, where the input prompt is compressed with 2k tokens. Moreover, as GemFilter focuses on the selected tokens more intensively than a full-context model by only processing selected tokens, with sufficient KV budget, there are some cases where it achieves even better accuracy than the full-context model. However, as the removal of tokens from the input prompt results in the complete loss of any information embedded in the discarded tokens, GemFilter struggles to maintain accuracy with lower KV budgets, such as 512, and even with 2k KV budgets when handling complex tasks. Thus, there are currently no KV cache compression techniques that can adequately preserve accuracy while simultaneously improving TTFT/throughput. This underscores the need for new methods that effectively balance these objectives.
3 Proposed FastKV
3.1 Overview of FastKV
The overview of FastKV is illustrated in Figure 2(e) and Algorithm 1. FastKV implements distinct KV cache compression strategies for the early and later layers of LLMs. To speedup the prefill stage, FastKV employs the TSP approach in the later layers. This approach is based on the insight that after the early layers have analyzed contextual information, the later layers tend to focus consistently on the same crucial tokens. The TSP approach, therefore, involves selectively propagating only a limited set of tokens (e.g. 2k tokens) identified by the TSP layer, which is strategically positioned in the middle of the LLM architecture. The TSP layer utilizes its attention map to pinpoint these critical tokens. Unlike the early layers, which propagate the entire set of tokens, the TSP approach in the later layers transmits only the selected tokens to subsequent layers. This targeted propagation means that the later layers process far fewer tokens, thus it can accelerate the processing speed in these layers. While the TSP approach results in a smaller set of KV data for the selected tokens, FastKV still applies the same KV cache compression techniques used in the early layers to optimize the compression ratio. In the early layers, FastKV evaluates the attention scores generated by each attention head to determine the importance of KV data. GQA compatible KV cache compression of FastKV selectively removes the KV data of tokens that receive low attention scores in an attention-group-wise manner. It is important to note that KV data compression occurs solely during the saving of KV data to establish KV caches for input prompt. Consequently, FastKV ensures that the entire set of KV data, including those of the removed tokens, is propagated to the subsequent layer until the TSP layer. This approach guarantees full context propagation to preserve essential contextual information from the input prompt. This dual strategy enables FastKV to achieve significant speed improvements without compromising accuracy, as it properly preserves the contextual information of input sequences.
3.2 Proposed TSP
TSP aims to enhance the efficiency of the prefill stage by propagating only a selected set of important tokens, similar to the approach used by GemFilter. However, in response to the insights obtained from the previous sections, TSP is designed to retain the necessary scope of token propagation to ensure that critical contextual information is preserved. The operation of TSP is governed by three key factors: the layer for token selection (TSP layer), identification of important tokens, and the number of tokens selected for propagation (TSP length). TSP begins by identifying TSP length important tokens from the full tokens at the TSP layer. The method for identifying important tokens in TSP builds upon Equation 1. The propagation of selected tokens across LLM layers involves transmitting the hidden states associated with those selected tokens. Consequently, TSP necessitates the assessment of token importance on a layer-by-layer basis. To facilitate this, TSP includes an additional step that aggregates the attention scores from all attention heads within the TSP layer as follows:
$S_{i}^{TSPlayer}=\frac{1}{H}\sum_{h=0}^{H-1}S_{i}^{TSPlayer,h}$ (2)
Here, $H$ denotes the number of attention head in each layer. From the TSP layer onward, TSP propagates only selected important tokens through the later layers. Unlike GemFilter, which reverts to the first layer of the LLM with selected tokens and recomputes the prefill stage, TSP directly propagates the selected tokens to subsequent layers without returning to the first layer. Meanwhile, although previous approaches to KV cache compression aimed at high accuracy by propagating the entire context throughout the prefill stage, the proposed TSP method discards the unimportant tokens after the TSP layer. Therefore, the implementation of TSP must be carefully designed based on a detailed analysis of the properties of token importance and its impact on the LLM output.
3.3 Impact of Important Tokens
Figure 3: Rate of sum of attention scores for the top 2k important tokens selected to the total attention score at each layer.
In the LLM, the attention mechanism is the critical component for extracting contextual information, while other parts of the model process input in a more token-isolated manner. Within the attention mechanism, each token generates a new embedding by integrating the embeddings of all preceding tokens, with this integration weighted by the attention scores. Consequently, tokens that receive low attention scores have only a marginal impact on the processing of contextual information. To ensure the removal of the unimportant tokens without significant impact on the LLM output, the important tokens must receive significantly higher attention scores.
We evaluate the actual impact of these important tokens by examining their attention scores relative to those of the remaining tokens, as shown in Figure 3. In this analysis, we calculate the sum of the attention scores assigned to 2k important tokens selected in each layer and compare this sum to the attention scores allocated to all 128k tokens. Despite important tokens comprising only 1.56% of the total number of tokens, they receive a majority of the attention scores. This is particularly notable in the later layers, where the proportion of attention scores attributed to these tokens consistently exceeds 50% and often approaches 80%. This result indicates that they are central to the LLM processing and the remaining tokens could be removed with a marginal impact on the LLM output.
3.4 Dynamics of Token Importance
Figure 4: Minimum match rate between indices of 2k important tokens selected at each layer and its subsequent layers.
Figure 5: Visualization of output logits by t-SNE for (left) LLaMA-3.1-8B-Instruct and (right) Mistral-Nemo-12B-Instruct.
Given that tokens discarded after the TSP layer are not propagated in subsequent layers, there must be consistency in identifying important tokens across layers to facilitate TSP. This section examines the dynamics of token importance across layers of LLMs. Specifically, we investigate whether tokens deemed important based on their high attention scores remain consistent across LLM layers. A significant overlap in the set of important tokens would suggest the potential for pre-selecting these tokens before propagating them to subsequent layers. To evaluate the consistency of important tokens across layers, we analyze the important tokens in each layer of the LLM and calculate the minimum match rate of the indices of these tokens with those in subsequent layers. The minimum match rate serves as an indicator of an information bottleneck.
Figure 4 presents our findings from this analysis on various LLMs, where each layer extracts 2k important tokens from 128k tokens. The minimum match rate in the first layer is below 10%, indicating low correlation between important tokens in the first and later layers. However, as we progress to subsequent layers, this match rate increases, surpassing 25% by the middle layer. This indicates that among the 2k tokens, at least 512 consistently appear as important across layers. Therefore, if the KV cache budget per layer is set to 512, it suffices to examine the 2k important tokens selected from the middle layer. The observed consistency in important token selection in later layers likely stems from the increasing robustness of contextual information extraction. Each layer in an LLM uses attention scores to integrate token information, progressively enhancing contextual understanding. As the model advances through its layers, contextual analysis becomes more refined, enabling a more precise identification of relevant tokens. Thus, later layers of LLMs exhibit a consistent recognition of important tokens, underscoring their refined understanding of token importance in the context of the overall sequence.
3.5 Overall Impact of TSP on LLM Output
Figure 6: LongBench results of LLaMA-3.1-8B-insturct with propagation of tokens selected at each layer.
Building on the insights from Section 3.3 and 3.4, we can expect that the removal of unimportant tokens identified from the middle layer will not significantly impact the performance of the LLM for two main reasons. First, only a small set of important tokens plays an important role in the attention mechanism and there is the consistency in the identification of important tokens in the middle layer and subsequent layers. We analyze real changes in the output logits of an LLM after selecting 2k tokens from each layer and discarding the unselected tokens before propagating them to subsequent layers. For this analysis, we use t-SNE (t-distributed Stochastic Neighbor Embedding) to project the final output logits of LLMs into a 2D space. Figure 5 demonstrates the distances between output logits obtained with full-context propagation and those obtained through propagation focusing exclusively on important tokens. This visual representation shows that removing tokens in the early layers causes significant deviations in output logits compared to full-context propagation. However, as expected, removing tokens from later layers preserves a high degree of similarity to the results from full-context models. This contributes to stable accuracy results on the LongBench benchmark, as illustrated in Figure 6. In light of this observation, we designate the middle layer of each LLM as the TSP layer for the remainder of this paper.
Moreover, in Figure 6, we compared the accuracy obtained with TSP and a GemFilter-like propagation. TSP consistently outperforms GemFilter in terms of accuracy. This superior performance can primarily be attributed to how TSP manages the variability in token importance across layers. Unlike GemFilter, which reverts to the first layer of the LLM with selected tokens and recomputes the prefill stage, TSP directly propagates the selected tokens to subsequent layers without reverting to the initial layer. This direct propagation allows TSP to better accommodate the variability in token importance. Furthermore, the large accuracy gap between TSP and GemFilter when tokens are selected in the initial layers highlights the importance of leveraging token information that has been processed with the full context before making important token selections.
3.6 TSP Length and KV Budget
The match rate of the important tokens is around 25% (Figure 4). This rate is sufficiently high to justify the use of TSP approach with an adequate TSP length (e.g. 2k). However, a short TSP length (e.g. 512) may lead to significant information loss. Therefore, TSP length, which defines the number of tokens selected for propagation, is strategically designed to operate independently of the KV cache budget. For instance, if the KV cache budget is set to 512 while the TSP length is 2048, TSP propagates 2048 tokens during its operation, but only the final 512 tokens are stored in the KV cache. This separation allows TSP to process a significantly larger set of tokens during propagation, enabling it to capture more contextual information while maintaining the KV cache within its predefined budget. As shown in Figure 7, a TSP length of 2048 has been found sufficient to preserve accuracy of the model under specific KV budget while maintaining efficiency in terms of TTFT.
@@ 3.6 TSP Length and KV Budget
Single-Doc QA Multi-Doc QA Summarization Few-shot Learning Coding
Method
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
LCC
RB-P
Avg.
LLaMA-3.1-8B-Inustruct, KV Budget = Full
Full KV 30.21 45.53 55.01 56.01 46.65 31.28 35.13 25.28 27.25 73.00 91.64 43.80 63.38 56.64 48.63
@@
SnapKV 30.09 41.62 53.89 54.77 45.12 31.22 27.37 24.16 24.86 71.00 91.90 42.43 61.53 52.49 46.60
AdaKV 29.07 40.16 52.44 53.90 43.05 31.10 25.75 24.39 24.85 69.00 92.34 42.05 63.43 55.32 46.20
HeadKV 30.17 44.03 54.29 54.71 46.10 31.58 31.14 24.46 26.66 73.00 91.72 42.17 62.89 55.71 47.76
GemFilter 18.52 22.15 37.73 46.88 32.56 24.52 27.05 17.75 23.25 60.50 89.49 40.16 27.35 31.02 35.64
FastKV 30.38 41.12 55.71 54.35 46.69 30.89 26.78 23.82 24.37 72.50 92.04 42.64 62.60 52.74 46.90

### v2

@@ A.1 FastKV Algorithm
1:$input$ $sequence$$\{I\},$$\#layers$$\{L\},$ $L_{TSP},$ $B_{TSP},$ $B_{KV}$
@@
5:     if $l\leq L_{TSP}$ then
6:         $X,Att_{l},K_{X},V_{X}\leftarrow layer_{l}(X)$triangleright $X:$ $hidden$ $states$ $before$ $TSP$
@@
8:         if $l==L_{TSP}$ then
9:              $x\leftarrow HiddenCompress(X,Att_{l},B_{TSP})$triangleright $x:$ $reduced$ $hidden$ $states$
@@
A.2 Comparison with GemFilter
We compare FastKV with GemFilter, a recent method that accelerates the prefill stage by pruning the input prompt based on token relevance. As shown in Figure 9, GemFilter selects a subset of tokens (e.g., 512) at an intermediate filter layer and then restarts the prefill computation using only these tokens. This two-stage process, consisting of preprocessing followed by recomputation, enables latency reduction in the prefill stage and effective KV cache compression under a fixed budget.
However, GemFilter’s strategy of pruning the input prompt leads to an irreversible loss of information from discarded tokens. Since only the selected tokens are used for recomputation, the hidden states of the removed tokens are not propagated to the later layers, fundamentally limiting the model’s ability to retain global context. As illustrated in Figure 10, this disrupts the full-sequence attention patterns essential for deep semantic understanding. As a result, GemFilter shows poor performance on tasks that require comprehensive context, such as code completion.
In contrast, FastKV mitigates this issue by introducing Token-Selective Propagation (TSP). Rather than pruning the input prompt, FastKV allows the early layers to process the full sequence and selects a subset of semantically important tokens (e.g., 2K) at a mid-layer (TSP layer) for propagation. The hidden states of all tokens are preserved and propagated until the TSP layer, ensuring that downstream layers receive semantically rich representations. This design enables FastKV to preserve contextual fidelity while still achieving significant latency and memory savings.
Figure 9: Comparison of processing flows between GemFilter and FastKV. GemFilter prunes the input prompt and recomputes later layers with only selected tokens, while FastKV propagates the full-context to a mid-layer before reducing the token set for further computation.
Figure 10: Visualization of attention computation. In GemFilter, discarded tokens are never represented in the attention output, leading to fragmented input understanding. FastKV retains all tokens through early layers and compresses only after contextual aggregation, preserving global semantics.
Appendix B Details on Evaluation
B.1 Models
We conduct our main evaluation using two open-source LLMs: LLaMA-3.1-8B-UltraLong-1M-Instruct and Mistral-Nemo-12B-Instruct. Additionally, we include LLaMA-3.1-8B-Instruct, the base model from which LLaMA-3.1-8B-UltraLong-1M-Instruct is further trained, for supplementary measurements. LLaMA-3.1-8B-Instruct supports a context window of 128K tokens.
The checkpoints for all models are publicly available at:
B.2 Datasets
In this section, we provide detailed information on the benchmark datasets used to evaluate FastKV’s performance against baseline methods.
LongBench. LongBench is a collection of benchmark datasets designed to assess the long-context understanding capabilities of LLMs. It includes tasks in both English and Chinese, spanning six categories, each consisting of multiple subtasks.
In this study, we evaluate FastKV on a subset of LongBench that includes only the English-language tasks and code-related benchmarks. Detailed information for each selected subtask is provided in Table 3.
Table 3: LongBench dataset details.
Task Type Task Abbreviation Eval metric Language #Sample
Single-doc QA NarrativeQA NrtvQA F1 EN 200
Qasper Qasper F1 EN 200
MultiFieldQA-en MF-en F1 EN 150
HotpotQA HotpotQA F1 EN 200
Multi-doc QA 2WikiMultihopQA 2WikiMultihopQA F1 EN 200
MuSiQue MuSiQue F1 EN 200
Summarization GovReport GovReport Rouge-L EN 200
QMSum QMSum Rouge-L EN 200
MultiNews MultiNews Rouge-L EN 200
Few shot TREC TREC Accuracy EN 200
TriviaQA TriviaQA F1 EN 200
SAMSum SAMSum Rouge-L EN 200
Code LCC LCC Edit Sim Python/C#/Java 500
RepoBench-P RB-P Edit Sim Python/Java 500
Synthetic PassageCount PCount Accuracy EN 200
PassageRetrieval-en PRe Accuracy EN 200
Needle-in-a-Haystack. Needle-in-a-Haystack is a benchmark designed to evaluate the in-context retrieval capabilities of LLMs under long-context settings. A statement, referred to as the needle, is randomly selected and inserted at a specific position within a long input context. The model is tasked with retrieving the needle content given the entire context. Evaluation is conducted across various context lengths and needle depths (insertion positions) to systematically measure retrieval performance under different levels of difficulty. In our experiments, retrieval performance is evaluated at context lengths ranging from 32K to 256K tokens, with measurements taken at 32K-token intervals.
Appendix C More Experimental Results
C.1 LongBench
We provide additional LongBench results for LLaMA-3.1-8B-UltraLong-1M-Instruct and Mistral-Nemo-12B-instruct with a KV budget of 1024 (Table 4), as well as a full breakdown of the LongBench results of an additional model, LLaMA-3.1-8B-Instruct (Table 5). LLaMA-3.1-8B-Instruct shares the same architecture as LLaMA-3.1-8B-UltraLong-1M-Instruct and serves as its base model, thus exhibiting similar performance trends to the two LLMs reported in Section 4.2.
Table 4: LongBench results comparison between baseline methods and FastKV with KV budget 1024.
@@ C.1 LongBench
Single-Document QA Multi-Document QA Summarization Few-shot Learning Code Synthetic
Method NrtvQA Qasper MF-en HotpotQA 2WikiMQA MuSiQue GovReport QMSum MultiNews TREC TriviaQA SAMSum LCC RB-P PCount PRe Avg.
LLaMA-3.1-8B-UltraLong-1M-Instruct, KV Budget = Full
Full KV 23.59 39.33 51.80 56.89 45.53 34.66 33.57 22.96 26.10 73.50 90.88 42.49 62.03 67.05 2.00 89.50 47.62
LLaMA-3.1-8B-UltraLong-1M-Instruct, KV Budget = 1024
SnapKV 23.82 37.46 51.63 56.66 45.65 34.55 29.13 22.78 25.31 73.50 91.09 41.65 62.08 65.94 2.00 89.00 47.02
AdaKV 23.46 37.49 51.54 56.20 44.98 34.91 29.19 22.76 25.31 73.50 91.06 42.37 62.18 67.40 2.00 88.50 47.05
HeadKV 23.60 37.97 52.03 55.99 45.14 34.92 31.15 23.44 26.12 73.50 90.88 41.28 62.61 67.15 2.00 89.00 47.30
GemFilter 17.68 29.73 52.20 53.25 49.03 28.34 29.73 20.46 24.46 67.50 88.84 38.43 43.08 45.89 1.00 74.00 41.48
FastKV 24.12 37.51 52.93 56.65 46.83 34.12 29.09 22.60 25.19 75.00 91.06 42.17 61.96 66.91 2.00 87.50 47.23
Mistral-Nemo-12B-Instruct, KV Budget = Full
Full KV 26.27 43.64 58.11 49.34 45.85 26.26 31.31 24.15 26.08 75.00 89.66 44.32 68.58 68.11 1.50 98.00 48.51
Mistral-Nemo-12B-Instruct, KV Budget = 1024
SnapKV 23.27 41.42 56.01 47.85 46.09 25.83 25.97 23.70 25.37 74.00 89.82 43.48 68.46 65.84 0.50 97.50 47.19
AdaKV 24.29 41.50 55.45 47.70 46.72 26.12 24.94 22.96 24.84 74.00 89.82 44.11 68.05 65.29 1.00 98.00 47.77
HeadKV 25.94 41.38 56.45 48.22 45.90 25.05 28.67 23.72 25.72 75.00 89.84 42.64 68.16 66.60 1.50 98.50 47.71
GemFilter 26.94 40.67 53.63 55.82 56.90 38.23 28.68 20.76 25.34 69.00 87.32 42.49 44.33 45.41 4.50 87.00 45.44
FastKV 26.26 42.37 57.71 49.40 46.88 25.58 26.17 23.08 25.29 75.00 89.72 43.56 70.13 67.39 0.50 99.00 48.00
Table 5: LLaMA-3.1-8B-Instruct LongBench results comparison between baseline methods and FastKV.
Single-Document QA Multi-Document QA Summarization Few-shot Learning Code Synthetic
Method NrtvQA Qasper MF-en HotpotQA 2WikiMQA MuSiQue GovReport QMSum MultiNews TREC TriviaQA SAMSum LCC RB-P PCount PRe Avg.
LLaMA-3.1-8B-Instruct, KV Budget = Full
Full KV 30.21 45.53 55.01 56.01 46.65 31.28 35.13 25.28 27.25 73.00 91.64 43.80 63.38 56.64 8.91 99.50 49.33
@@
SnapKV 30.09 41.62 53.89 54.77 45.12 31.22 27.37 24.16 24.86 71.00 91.90 42.43 61.53 52.49 8.62 99.50 47.54
AdaKV 30.16 40.11 54.13 54.27 44.73 31.19 27.08 23.82 24.39 71.00 91.90 41.64 60.54 53.15 8.12 99.50 47.23
HeadKV 30.17 44.03 54.29 54.71 46.10 31.58 31.14 24.46 26.66 73.00 91.72 42.17 62.89 55.71 8.33 99.50 48.53
GemFilter 18.52 22.15 37.73 46.88 32.56 24.52 27.05 17.75 23.25 60.50 89.49 40.16 27.35 31.02 3.53 52.00 34.65
FastKV 30.28 41.08 55.50 54.35 46.69 30.89 26.81 24.01 24.38 72.50 92.04 42.67 63.99 54.94 7.06 99.50 47.92
LLaMA-3.1-8B-Instruct, KV Budget = 1024
SnapKV 31.23 42.52 53.96 55.48 45.43 31.50 29.54 24.78 26.17 70.50 91.73 42.52 62.25 54.72 7.96 99.50 48.11
AdaKV 30.83 44.33 54.54 55.86 45.39 31.96 28.70 24.67 26.10 71.50 91.48 42.40 61.71 54.81 8.06 99.50 48.24
HeadKV 30.55 44.66 54.69 55.47 46.20 31.63 32.88 25.14 27.10 73.00 91.57 42.88 63.76 56.06 8.36 99.50 48.97
GemFilter 19.34 34.76 46.58 46.82 42.83 27.51 29.89 18.96 25.68 63.00 90.70 42.50 38.09 35.14 3.82 86.50 40.76
FastKV 30.25 44.01 54.72 54.78 46.56 30.59 29.38 24.54 26.23 73.50 91.76 42.58 63.99 57.32 6.86 99.50 48.54